Using identities, evaluate. 
998^2 ????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

998^2= (1000 - 2)2
= 1000^2- 4000 + 2^2
= 1000000 - 4000 + 4
= 996004

Thanks